OpenStudy (anonymous):

what about 9g^2+21gs+10s^2?

OpenStudy (dumbcow):

To factor multiply outside coefficients 9*10 = 90 look for factors of 90 that add up to 21 2*45 3*30 6*15 9*10 6+15=21 (9g +6s)(9g+15s) reduce (3g+2s)(3g+5s)